OpenStudy (anonymous):

How many hydrogen atoms in a baby aspirin tablet which has 81mg of aspirin? (aspirin is C9H804)

OpenStudy (anonymous):

allisante r u in chemistry 1st semester?

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

i started with 81mg aspirin, divided by 1000mg to get grams, then divided by 180.154 grams (the molar mass of aspirin), then multiplied by 8 moles of hydrogen, then multiplied by avogadros number. Im not sure if that is correct though...

OpenStudy (anonymous):

allisante

OpenStudy (anonymous):

i think the answer is 14590.943505

OpenStudy (aaronq):

yep. it's correct.

OpenStudy (anonymous):

samsterz, how did you get the answer though? aaronq, which answer is correct?

OpenStudy (anonymous):

how did i get the correct answer? did some math

OpenStudy (anonymous):

well thats the whole problem im having. i dont want someone to just hand me the answer. i want to KNOW how to do it so that i can do it myself too. i need to figure out how to set up the equation.

OpenStudy (doc.brown):

Grab your periodic table and look up the amu for each element. Add together the weights per mole, for example, the amu for Carbon (C) is 12.011g/mol \[9(C) + 8(H) + 4(O) = total\frac{g}{mol}\] convert mg to g to mol to atoms \[81mg \times \frac{g}{1000mg} \times \frac{mol}{total g} \times \frac{6.022\times10^{23}atoms}{mol}\]

OpenStudy (anonymous):

Thank you! so, for the mol/totalg part, i write 8mol/180.154g?

OpenStudy (doc.brown):

\[9C+8H+4O\]\[9(12.011\frac{g}{mol}) + 8(1.008\frac{g}{mol}) + 4(15.999\frac{g}{mol})\]\[=18.159\frac{g}{mol}\]We're saying here that 18.159g of our molecule is the same mass as 1 mole of our molecule so we can also say\[\frac{1 mol}{18.159g}\]

OpenStudy (doc.brown):

Here it is in full blown scientific notation, just for practice \[8.1\times10^{1}mg\times\frac{g}{10^{3}mg}\times\frac{mol}{1.80159\times10^{2}}\times\frac{6.022\times10^{23}atoms}{mol}\]\[8.1\times6.022\div1.80159\]\[=27.075=2.7\times10^{1}\]add up the exponents \[10^{1+1-3-2+23}=10^{20}\]therefore you have\[2.7\times10^{20}atoms\]to two significant figures because you have two sig.fig. in 81mg

OpenStudy (anonymous):

thank you!